an orange grower has 400 crates or fruit ready and will have 20 more for each day the grower waits to sell. the present price is $60 per crate and will drop $2 per crate per day. in how many days should the grower ship the crop to maximize income?
Found 2 solutions by nerdybill, stanbon:
Answer by nerdybill(7384)

.
Let x = number of days until shipped
then
f(x) = (60-2n)(400 + 20n)
where f(x) is the income
.
f(x) = (60-2n)(400 + 20n)
f(x) = 24000+1200n-800n-40n^2
f(x) = -40n^2+400n+24000
.
Since the coefficient for the n^2 term is -40 (negative), we know that it is a parabola the opens downward. So finding the axis of symmetry will give us the max.
.
axis of symmetry = -b/2a = -400/2(-40) = -400/-80 = 5
.
Answer: 5 days

Answer by stanbon(75887)

an orange grower has 400 crates of fruit ready and will have 20 more for each day the grower waits to sell. the present price is $60 per crate and will drop $2 per crate per day. in how many days should the grower ship the crop to maximize income?
---------------------------------------
Revenu = price * quantity
R(x) = (60 - 2x)(400 + 20x)
R(x) = 2(30-x)*20(20+x)
R(x) = -40(x-30)(x+20)
R(x) = -40(x^2 - 10x - 600)
--------------------------------
Max of min occurs when x = -b/2a = 10/2 = 5
Maximum revenue will occur on day 5
============================================
Cheers,
Stan H.